If you buy a call option on a Birr 200,000 bond futures contract with an exercise price of 220 and the
price of the Treasury bond is 222 at expiration, is the contract in the money, out of the money, or at
the money? What is your profit or loss on the contract if the premium was Birr 3000?
1
Expert's answer
2020-08-17T19:27:01-0400

The call option is in the money.

200000/220= 909

222-222=2

909*2=1818

3000-1818=1182 Loss


Explanation :

If a call option's strike price or exercise price is lower than the market price at the exercise date, therefore that contract is called in the money. In the problem, the strike price is 220 while the market price is 222. Strike price is greater.


For the answer regarding the profit or loss, the assumption is that birr200000 of future contract will have 909 bonds with value of 220. the change in price of 222 means that there is a change in the prices. if the situation occurs the company will have a gain of 2 pesos per bond. The problem is the option premium of three thousand. these premium is a cost in part of the entity therefore the company incurred 1182 loss.